# Recursive numeral systems are highly regular and easy to process

### v1

@@ Abstract
Previous work has argued that recursive numeral systems optimise the trade-off between lexicon size and average morphosyntatic complexity (Denić and Szymanik, 2024). However, showing that only natural-language-like systems optimise this tradeoff has proven elusive, and the existing solution has relied on ad-hoc constraints to rule out unnatural systems (Yang and Regier, 2025). Here, we argue that this issue arises because the proposed trade-off has neglected regularity, a crucial aspect of complexity central to human grammars in general. Drawing on the Minimum Description Length (MDL) approach, we propose that recursive numeral systems are better viewed as efficient with regard to their regularity and processing complexity. We show that our MDL-based measures of regularity and processing complexity better capture the key differences between attested, natural systems and unattested but possible ones, including “optimal” recursive numeral systems from previous work, and that the ad-hoc constraints from previous literature naturally follow from regularity. Our approach highlights the need to incorporate regularity across sets of forms in studies that attempt to measure and explain optimality in language.
@@ 1 Introduction
One key question in the study of language is why languages are the way they are; what cognitive and communicative constraints shape linguistic systems? Recent research suggests that linguistic systems are shaped by efficient communication, such that natural systems optimise the trade-off between simplicity and informativeness in various domains (Gibson et al., 2019), from kinship terms (Kemp and Regier, 2012), colour terms (Regier et al., 2015; Zaslavsky et al., 2019), grammatical categories (Mollica et al., 2021), indefinite pronouns (Denić et al., 2022), boolean connectives (Uegaki, 2022) to spatial demonstratives (Chen et al., 2023). For example, colour-naming systems tend to either partition the space of colour referents into few large, connected categories (e.g., dark vs. light colours), or into many small specific categories (e.g., red, green, blue, yellow, brown, …); the former systems are simple but do not allow precise communication about colour, the terms in the latter systems are highly informative but the systems are complex (Regier et al., 2015).
Xu et al. (2020) investigated this trade-off in the domain of numeral systems. They found that while numeral systems with limited precision (e.g., a few vs. many) followed the same trade-off, recursive numeral systems like the English decimal system (one, two, … ninety, ninety-one, …) are maximally informative in that any integer can be precisely expressed. Since informativeness is fixed, Denić and Szymanik (2024) (henceforth, D&S) argued that the relevant trade-off shaping natural recursive numeral systems is instead between lexicon size
and the average length of the numerals. However, as reviewed in Section 2, the optimal numeral systems under this trade-off included not only human-like systems, but also unnatural, highly irregular ones.
Yang and Regier (2025) (Y&R) attempted to resolve this with ad-hoc constraints that artificially restrict the space of possible systems to be human-like.
@@ 2 Prior work on efficiency of recursive numeral systems
To show that natural recursive systems optimise this trade-off, D&S compared the numeral systems of natural languages to theoretically possible systems over the number range 1991-99. They generated these possible recursive numeral systems using Hurford’s grammar (Hurford, 1975, 2007):
@@ 19 $10+9$ 25625-6
Y&R attempted to solve these issues by introducing ad-hoc constraints over the space of possible numeral systems, resolving the unorderliness issue of D&S by restricting the numeral system space to be more human-like.
They added an extra rule for suppletives, and four additional constraints, which specify which multipliers to use for which number ranges, the maximum number that can be added or subtracted from Phrase{Phrase} within a given number range, as well as exception rules.
They further require that the digits in $D$ run sequentially from $1$ to a number between 1201-20. They also only allowed for the smallest multiplier to be used to construct numbers smaller than or equal to the next largest multplier in $M$ (Other examples can be found in Yang and Regier, 2025).
Crucially, Y&R still maintain that lexicon size and morphosyntactic complexity are the relevant measures.
In this sense, D&S’s optimal languages are still considered as efficient as Y&R ones, even though they cannot be generated under the grammar constraints used by Y&R.
@@ 3.1 Irregularity
One way to measure regularity is in terms of compressibility. Regular systems are highly compressible, such that one could write a small set of general rules for the construction of numerals. For instance, recursive numerals in Karo Batak (Woollams, 1996) are extremely regular. For numerals from 1 to 99, it only requires the rules $Num\rightarrow{}D$ (for 1-9), and $Num\rightarrow{}D*10+D$ (for 10-99).
English numerals are slightly less regular, i.e. featuring suppletive numerals ($11$, $12$) that require separate rules. We can thus measure the (ir)regularity of a system by finding the length of the most general/shortest grammar for numerals. However, as inferring the shortest recursive grammar is challenging (Charikar et al., 2005), we simplify the problem by restricting ourselves to the number range 1-99, following D&S and Y&R, and treating the system as a finite language.
We inferred a minimal partial deterministic finite automaton (DFA) that generates all and only valid numerals to represent a numeral grammar using Mihov and Schulz’s (2019) algorithm. [^1] We use the Python package automata-lib (https://pypi.org/project/automata-lib/) for inference. Intuitively, a regular system has a lot of repeated parts, so the number of states and transitions of the DFA will be small (cf. Figure 1).
@@
Our measure of irregularity is thus the representational complexity of the numeral system, i.e., the size of the partial DFA. Adapting the scheme of Brighton (2003), the irregularity $L(G)$ of a recursive numeral grammar $G$, in bits, is:
@@ 3.2 Processing complexity
The DFA can be thought of as a description language used to encode the data (a set of numerals). The more complex the description language (the larger the DFA), the easier it is to generate or parse a numeral. Drawing on the Minimum Description Length principle (Li et al., 2008), Brighton (2003) shows that this quantity could be calculated from the average information required to determine the unique parsing path through the DFA for a given form-meaning pairing. In our case, this pairing simply corresponds to a numeral (the form) as one morpheme refers to only one meaning (e.g., one in twenty one refers only to ‘1’). For instance, in the DFA shown in Figure 1, parsing 2102*10 (20) requires going through 4 transitions connecting the states $\lambda{}$ (initial state), $D$, DD* and D10D*10, and one would need $\underbrace{log_{2}(9)}_{\lambda}+\underbrace{log_{2}(1)+1}_{D}+\underbrace{log_{2}(1)}_{D*}+\underbrace{log_{2}(1)+1}_{D*10}$ bits to encode or generate this numeral as a path through the DFA, with each accepting state requiring an extra bit to encode the accepting status.
Averaging over these bits for all numbers gives the processing complexity of generating or parsing a set of numerals using a given DFA. Again adapting Brighton’s (2003) scheme, the processing complexity $L(N\mid{}G)$ of recursive numeral grammar $G$ for a set of numbers $N$ is given by:
@@
where $P(n_{i})$ is the prior/weight for $n_{i}\in N$, $S_{i}$ is the list of states through which the parsing proceeds for the numeral for $n_{i}$, $Z_{ij}$ is the set of outward transitions from the $j$th state in the parsing of the numeral for $n_{i}$, and $F$ is an indicator function evaluating to 1 if $s_{ij}\in S_{i}$ is an accepting state. D&S and Y&R assume that the prior is $P(n_{i})\propto n_{i}^{-2}$, taken to reflect the communicative need of numbers (Dehaene and Mehler, 1992; Piantadosi, 2016), i.e. there is high communicative need / frequent use of small numbers, and the need to express higher numbers declines rapidly. We follow this assumption in Section 4 but challenge it in Section 5.
@@ 4 Applying the measures
We first ask if natural recursive numerals are particularly regular and of low processing complexity, relative to the entire space of possible systems. We compute our measures on two datasets: (1) the recursive numeral systems from 128 natural languages annotated at the morpheme level from D&S, [^2] https://osf.io/cmuq4 and (2) a sample of 10,000 baseline artificial systems generated with Hurford’s grammar. To obtain the baseline sample, we adapt the implementation of the genetic algorithm in D&S. When constructing artificial numeral systems, in cases where a grammar allows a number to be expressed as several possible numerals, D&S always chose the shortest numeral for each number (e.g. given $D=\{1,2\},M=\{5,10,12\},$ the possible numerals for 12 include the shortest $12$, along with 1121*12, $10+2$, 110+21*10+2 and 25+22*5+2). This is partly the source of unorderliness: favouring the shortest possible numeral will result in a system that does not reuse structures in smaller numerals in bigger ones (e.g., using $12$ instead of $10+2$ when the numeral for 11 is $10+1$). To alleviate this, we allow numerals of any length and any combination of digits, multipliers and combinators (+,,+,*,-). For computational tractability, however, we limit the max search depth to 5, meaning that a numeral has at most 5 number morphemes and 4 arithmetic combinators, and use only digits and multipliers that occur in at least one natural language. We draw languages in 100 batches of 100 language types. We determine the type for each batch as follows. We randomly select 3 to 12 digits and 1 to 3 multipliers. We always allow ++ and * as combinators, with a 20% chance of also having the subtraction combinator -, which is less common among the world’s languages (Comrie, 2022). We resample if the resulting set of digits, multipliers and combinators cannot express all numbers between 1 and 99. Otherwise, we use the selected digits, multipliers and combinators to generate 100 sets of numerals, each covering the numbers 1-99, where those samples can differ in how a given number is expressed (cf. the multiple possible numerals for 12 above).
@@ 5 The role of the prior
Previous work has argued that recursive numeral systems are subjected to a power law prior distribution $P(n)\propto n^{-2}$, which reflects communicative need over numbers (Dehaene and Mehler, 1992; Xu et al., 2020; Denić and Szymanik, 2024; Yang and Regier, 2025). This presents a very strong bias towards the smallest numbers, as when calculating either average morphosyntactic or processing complexity most numerals will have minimal influence. More specifically, when considering the numeral range 1991-99, more than $75\%$ of the probability mass is placed on numbers $1$ and $2$ alone, and more than $90\%$ is placed on numbers 161-6. This is why, for example, the optimal system in Section 4.1 with incredibly complex large numerals (e.g. a system where the number 46 expressed as the numeral (((22+1)2+1)2+1)2(((2*2+1)*2+1)*2+1)*2) is considered easier process than natural systems: these numerals contribute negligibly to the final result when considering the power law prior.
@@ 6 Discussion
We have argued that prior literature on the efficiency of recursive numeral systems (Denić and Szymanik, 2024; Yang and Regier, 2025) has not satisfactorily accounted for the key characteristics of human-like systems, namely their regularity and low processing complexity. In this paper, we have shown that human recursive numeral systems are more regular and easier to process than other theoretically possible systems, even when controlling for the set of number, morphemes and numeral lengths. We are skeptical that the relevant prior is the power law communicative need distribution (Dehaene and Mehler, 1992; Piantadosi, 2016) as assumed in prior work (Xu et al., 2020; Denić and Szymanik, 2024; Yang and Regier, 2025). We show that even if the relevant prior was uniform, our measures would still clearly separate human-like and other possible recursive numeral systems.
@@
One issue that we have not considered here is the role of cognitive cost of encoding certain meanings or relations. For example, systems using highly unnatural numerals like (((22+1)2+1)2+1)2(((2*2+1)*2+1)*2+1)*2 (Section 4.1) are currently considered efficient. We have shown that this might be because of the power law prior assumption (Section 5), but the absence of this kind of numeral could also be due to the high cost of computation involved in the multiply nested multiplications, which are harder to learn in numeral systems than operations like additions (Holt and Barner, 2025; Silvi et al., 2025), while our measure of processing complexity assumes that all transition symbols are equally costly. Similarly, expressing numerals with multiple nested parentheses would also introduce ambiguity in how they are interpreted (e.g., 210+52*10+5 could be either (210)+5(2*10)+5 or 2(10+5)2*(10+5)).
@@
Another issue concerns the granularity of the annotated data from D&S. We notice that the annotation was not sensitive to the order of morphemes in each numeral, and they would always annotate the numerals according to Hurford’s left-headed grammar. For example, English -teens were annotated as $10+D$ instead of the real order $D+10$ (e.g., four-teen should be $4+10$). The same is true for multiplications; the product of $D$ and $M$ is annotated as DMD*M. For example, ìri īse ‘50’ in Igbo is 10510*5 (Emenanjo, 1987) but was annotated 5105*10. This presents no problem for their measures, which are not sensitive to morpheme order. However, our measures are, and this annotation results in overestimated regularity, since, for instance, the orders of $M$ and $D$ in natural languages often change depending on number ranges, reducing regularity. We have begun the process of reannotating their data, in order to re-run the analyses here taking this order variation into account. Early results suggest that the order-sensitive reannotation does not change our qualitative results, at least with regard to the efficiency of recursive numeral systems relative to other theoretically possible, including D&S- and Y&R-optimal systems. This is likely because natural systems maintain a high degree of regularity in order over large number ranges (e.g., only use $M+D$ over the teens range, as in English).
@@
Unlike most work in the efficient communication literature, which only considers the mapping between form and meaning (cf. Kemp et al., 2018), work on recursive numeral systems like ours, D&S and Y&R focuses on the forms themselves. Our work specifically examines the relationship among forms, i.e., the regularity of the system, and shows that it is a vital property of efficient, human-like linguistic systems. This suggests, for example, that simplicity might be underestimated in work that treats compositional signals as equally complex as unanalysable ones (e.g., in grammatical categories like person systems; Zaslavsky et al., 2021). Moreover, our re-examination of D&S and Y&R also suggests that it is not sufficient simply to find a set of measures under which natural numeral systems could be shown to be optimal. To actually pinpoint why certain systems are human-like, one would want to identify the properties, like regularity and processing complexity, which explain the natural preferences of human-like systems over other possible ones, and which have plausible cognitive bases.
@@
One clear extension of our work here is an investigation into whether natural recursive numeral systems are on the Pareto frontier of the irregularity/processing complexity trade-off. This is particularly challenging, because the space of all possible recursive numeral systems is extremely large, even when we only consider Hurford-grammar-conforming systems and limit the maximum search depth. Unlike D&S and Y&R, we would also like to avoid constraining the space by simply favouring shortest numerals or adding ad-hoc constraints. One potential solution is to adopt Y&R’s enhanced grammars as representations while dropping their ad-hoc implementational constraints, allowing us to consider a much larger solution space while still keeping compact recursive numeral representations. Another complementary path would be to provide a mechanistic explanation of how these systems have been shaped over time via
general learning mechanisms such as reinforcement learning, similar to Carlsson et al. (2021) and Silvi et al. (2025), and also via behavioural experiments.

### v2

@@ Abstract
Previous work has argued that recursive numeral systems optimise the trade-off between lexicon size and average morphosyntatic complexity (denic2024recursive). However, showing that only natural-language-like systems optimise this tradeoff has proven elusive, and the existing solution has relied on ad-hoc constraints to rule out unnatural systems (yang2025re). Here, we argue that this issue arises because the proposed trade-off has neglected regularity, a crucial aspect of complexity central to human grammars in general. Drawing on the Minimum Description Length (MDL) approach, we propose that recursive numeral systems are better viewed as efficient with regard to their regularity and processing complexity. We show that our MDL-based measures of regularity and processing complexity better capture the key differences between attested, natural systems and unattested but possible ones, including “optimal” recursive numeral systems from previous work, and that the ad-hoc constraints from previous literature naturally follow from regularity. Our approach highlights the need to incorporate regularity across sets of forms in studies that attempt to measure and explain optimality in language.
@@ 1 Introduction
One key question in the study of language is why languages are the way they are; what cognitive and communicative constraints shape linguistic systems? Recent research suggests that linguistic systems are shaped by efficient communication, such that natural systems optimise the trade-off between simplicity and informativeness in various domains (gibson2019efficiency), from kinship terms (kemp2012kinship), colour terms (regier2015word; zaslavsky2019col), grammatical categories (mollica2021forms), indefinite pronouns (denicIndefinitePronounsOptimize2022a), boolean connectives (uegaki_informativenesscomplexity_2022) to spatial demonstratives (chen2023information). For example, colour-naming systems tend to either partition the space of colour referents into few large, connected categories (e.g., dark vs. light colours), or into many small specific categories (e.g., red, green, blue, yellow, brown, …); the former systems are simple but do not allow precise communication about colour, the terms in the latter systems are highly informative but the systems are complex (regier2015word).
xu2020numeral investigated this trade-off in the domain of numeral systems. They found that while numeral systems with limited precision (e.g., a few vs. many) followed the same trade-off, recursive numeral systems like the English decimal system (one, two, … ninety, ninety-one, …) are maximally informative in that any integer can be precisely expressed. Since informativeness is fixed, denic2024recursive (henceforth, D&S) argued that the relevant trade-off shaping natural recursive numeral systems is instead between lexicon size and the average length of the numerals. However, as reviewed in Section 2, the optimal numeral systems under this trade-off included not only human-like systems, but also unnatural, highly irregular ones. yang2025re (Y&R) attempted to resolve this with ad-hoc constraints that artificially restrict the space of possible systems to be human-like.
@@ 2 Prior work on efficiency of recursive numeral systems
To show that natural recursive systems optimise this trade-off, D&S compared the numeral systems of natural languages to theoretically possible systems over the number range 1991-99. They generated these possible recursive numeral systems using Hurford’s grammar (Hurford1975TheLT; HURFORD2007773):
@@ 19 $10+9$ 25625-6
Y&R attempted to solve these issues by introducing ad-hoc constraints over the space of possible numeral systems, resolving the unorderliness issue of D&S by restricting the numeral system space to be more human-like. They added an extra rule for suppletives, and four additional constraints, which specify which multipliers to use for which number ranges, the maximum number that can be added or subtracted from Phrase{Phrase} within a given number range, as well as exception rules. They further require that the digits in $D$ run sequentially from $1$ to a number between 1201-20. They also only allowed for the smallest multiplier to be used to construct numbers smaller than or equal to the next largest multplier in $M$ (Other examples can be found in yang2025re).
Crucially, Y&R still maintain that lexicon size and morphosyntactic complexity are the relevant measures. In this sense, D&S’s optimal languages are still considered as efficient as Y&R ones, even though they cannot be generated under the grammar constraints used by Y&R.
@@ 3.1 Irregularity
One way to measure regularity is in terms of compressibility. Regular systems are highly compressible, such that one could write a small set of general rules for the construction of numerals. For instance, recursive numerals in Karo Batak (woollamsGrammarKaroBatak1996) are extremely regular. For numerals from 1 to 99, it only requires the rules $Num\rightarrow{}D$ (for 1-9), and $Num\rightarrow{}D*10+D$ (for 10-99).
English numerals are slightly less regular, i.e. featuring suppletive numerals ($11$, $12$) that require separate rules. We can thus measure the (ir)regularity of a system by finding the length of the most general/shortest grammar for numerals. However, as inferring the shortest recursive grammar is challenging (charikarSmallestGrammarProblem2005), we simplify the problem by restricting ourselves to the number range 1-99, following D&S and Y&R, and treating the system as a finite language.
We inferred a minimal partial deterministic finite automaton (DFA) that generates all and only valid numerals to represent a numeral grammar using mihovFiniteStateTechniquesAutomata2019’s (mihovFiniteStateTechniquesAutomata2019) algorithm. [^1] We use the Python package automata-lib (https://pypi.org/project/automata-lib/) for inference. Intuitively, a regular system has a lot of repeated parts, so the number of states and transitions of the DFA will be small (cf. Figure 1).
@@
Our measure of irregularity is thus the representational complexity of the numeral system, i.e., the size of the partial DFA. Adapting the scheme of brighton2003simplicity, the irregularity $L(G)$ of a recursive numeral grammar $G$, in bits, is:
@@ 3.2 Processing complexity
The DFA can be thought of as a description language used to encode the data (a set of numerals). The more complex the description language (the larger the DFA), the easier it is to generate or parse a numeral. Drawing on the Minimum Description Length principle (li2008introduction), brighton2003simplicity shows that this quantity could be calculated from the average information required to determine the unique parsing path through the DFA for a given form-meaning pairing. In our case, this pairing simply corresponds to a numeral (the form) as one morpheme refers to only one meaning (e.g., one in twenty one refers only to ‘1’). For instance, in the DFA shown in Figure 1, parsing 2102*10 (20) requires going through 4 transitions connecting the states $\lambda{}$ (initial state), $D$, DD* and D10D*10, and one would need $\underbrace{log_{2}(9)}_{\lambda}+\underbrace{log_{2}(1)+1}_{D}+\underbrace{log_{2}(1)}_{D*}+\underbrace{log_{2}(1)+1}_{D*10}$ bits to encode or generate this numeral as a path through the DFA, with each accepting state requiring an extra bit to encode the accepting status.
Averaging over these bits for all numbers gives the processing complexity of generating or parsing a set of numerals using a given DFA. Again adapting brighton2003simplicity’s (brighton2003simplicity) scheme, the processing complexity $L(N\mid{}G)$ of recursive numeral grammar $G$ for a set of numbers $N$ is given by:
@@
where $P(n_{i})$ is the prior/weight for $n_{i}\in N$, $S_{i}$ is the list of states through which the parsing proceeds for the numeral for $n_{i}$, $Z_{ij}$ is the set of outward transitions from the $j$th state in the parsing of the numeral for $n_{i}$, and $F$ is an indicator function evaluating to 1 if $s_{ij}\in S_{i}$ is an accepting state. D&S and Y&R assume that the prior is $P(n_{i})\propto n_{i}^{-2}$, taken to reflect the communicative need of numbers (dehaene1992cross; piantadosi2016rational), i.e. there is high communicative need / frequent use of small numbers, and the need to express higher numbers declines rapidly. We follow this assumption in Section 4 but challenge it in Section 5.
@@ 4 Applying the measures
We first ask if natural recursive numerals are particularly regular and of low processing complexity, relative to the entire space of possible systems. We compute our measures on two datasets: (1) the recursive numeral systems from 128 natural languages annotated at the morpheme level from D&S, [^2] https://osf.io/cmuq4 and (2) a sample of 10,000 baseline artificial systems generated with Hurford’s grammar. To obtain the baseline sample, we adapt the implementation of the genetic algorithm in D&S. When constructing artificial numeral systems, in cases where a grammar allows a number to be expressed as several possible numerals, D&S always chose the shortest numeral for each number (e.g. given $D=\{1,2\},M=\{5,10,12\},$ the possible numerals for 12 include the shortest $12$, along with 1121*12, $10+2$, 110+21*10+2 and 25+22*5+2). This is partly the source of unorderliness: favouring the shortest possible numeral will result in a system that does not reuse structures in smaller numerals in bigger ones (e.g., using $12$ instead of $10+2$ when the numeral for 11 is $10+1$). To alleviate this, we allow numerals of any length and any combination of digits, multipliers and combinators (+,,+,*,-). For computational tractability, however, we limit the max search depth to 5, meaning that a numeral has at most 5 number morphemes and 4 arithmetic combinators, and use only digits and multipliers that occur in at least one natural language. We draw languages in 100 batches of 100 language types. We determine the type for each batch as follows. We randomly select 3 to 12 digits and 1 to 3 multipliers. We always allow ++ and * as combinators, with a 20% chance of also having the subtraction combinator -, which is less common among the world’s languages (comrie2022arithmetic). We resample if the resulting set of digits, multipliers and combinators cannot express all numbers between 1 and 99. Otherwise, we use the selected digits, multipliers and combinators to generate 100 sets of numerals, each covering the numbers 1-99, where those samples can differ in how a given number is expressed (cf. the multiple possible numerals for 12 above).
@@ 5 The role of the prior
Previous work has argued that recursive numeral systems are subjected to a power law prior distribution $P(n)\propto n^{-2}$, which reflects communicative need over numbers (dehaene1992cross; xu2020numeral; denic2024recursive; yang2025re). This presents a very strong bias towards the smallest numbers, as when calculating either average morphosyntactic or processing complexity most numerals will have minimal influence. More specifically, when considering the numeral range 1991-99, more than $75\%$ of the probability mass is placed on numbers $1$ and $2$ alone, and more than $90\%$ is placed on numbers 161-6. This is why, for example, the optimal system in Section 4.1 with incredibly complex large numerals (e.g. a system where the number 46 expressed as the numeral (((22+1)2+1)2+1)2(((2*2+1)*2+1)*2+1)*2) is considered easier process than natural systems: these numerals contribute negligibly to the final result when considering the power law prior.
@@ 6 Discussion
We have argued that prior literature on the efficiency of recursive numeral systems (denic2024recursive; yang2025re) has not satisfactorily accounted for the key characteristics of human-like systems, namely their regularity and low processing complexity. In this paper, we have shown that human recursive numeral systems are more regular and easier to process than other theoretically possible systems, even when controlling for the set of number, morphemes and numeral lengths. We are skeptical that the relevant prior is the power law communicative need distribution (dehaene1992cross; piantadosi2016rational) as assumed in prior work (xu2020numeral; denic2024recursive; yang2025re). We show that even if the relevant prior was uniform, our measures would still clearly separate human-like and other possible recursive numeral systems.
@@
One issue that we have not considered here is the role of cognitive cost of encoding certain meanings or relations. For example, systems using highly unnatural numerals like (((22+1)2+1)2+1)2(((2*2+1)*2+1)*2+1)*2 (Section 4.1) are currently considered efficient. We have shown that this might be because of the power law prior assumption (Section 5), but the absence of this kind of numeral could also be due to the high cost of computation involved in the multiply nested multiplications, which are harder to learn in numeral systems than operations like additions (holt2025learning; silvi2025le), while our measure of processing complexity assumes that all transition symbols are equally costly. Similarly, expressing numerals with multiple nested parentheses would also introduce ambiguity in how they are interpreted (e.g., 210+52*10+5 could be either (210)+5(2*10)+5 or 2(10+5)2*(10+5)).
@@
Another issue concerns the granularity of the annotated data from D&S. We notice that the annotation was not sensitive to the order of morphemes in each numeral, and they would always annotate the numerals according to Hurford’s left-headed grammar. For example, English -teens were annotated as $10+D$ instead of the real order $D+10$ (e.g., four-teen should be $4+10$). The same is true for multiplications; the product of $D$ and $M$ is annotated as DMD*M. For example, ìri īse ‘50’ in Igbo is 10510*5 (emenanjo1987elements) but was annotated 5105*10. This presents no problem for their measures, which are not sensitive to morpheme order. However, our measures are, and this annotation results in overestimated regularity, since, for instance, the orders of $M$ and $D$ in natural languages often change depending on number ranges, reducing regularity. We have begun the process of reannotating their data, in order to re-run the analyses here taking this order variation into account. Early results suggest that the order-sensitive reannotation does not change our qualitative results, at least with regard to the efficiency of recursive numeral systems relative to other theoretically possible, including D&S- and Y&R-optimal systems. This is likely because natural systems maintain a high degree of regularity in order over large number ranges (e.g., only use $M+D$ over the teens range, as in English).
@@
Unlike most work in the efficient communication literature, which only considers the mapping between form and meaning (cf. kemp2018semantic), work on recursive numeral systems like ours, D&S and Y&R focuses on the forms themselves. Our work specifically examines the relationship among forms, i.e., the regularity of the system, and shows that it is a vital property of efficient, human-like linguistic systems. This suggests, for example, that simplicity might be underestimated in work that treats compositional signals as equally complex as unanalysable ones (e.g., in grammatical categories like person systems; Zaslavsky2021person). Moreover, our re-examination of D&S and Y&R also suggests that it is not sufficient simply to find a set of measures under which natural numeral systems could be shown to be optimal. To actually pinpoint why certain systems are human-like, one would want to identify the properties, like regularity and processing complexity, which explain the natural preferences of human-like systems over other possible ones, and which have plausible cognitive bases.
@@
One clear extension of our work here is an investigation into whether natural recursive numeral systems are on the Pareto frontier of the irregularity/processing complexity trade-off. This is particularly challenging, because the space of all possible recursive numeral systems is extremely large, even when we only consider Hurford-grammar-conforming systems and limit the maximum search depth. Unlike D&S and Y&R, we would also like to avoid constraining the space by simply favouring shortest numerals or adding ad-hoc constraints. One potential solution is to adopt Y&R’s enhanced grammars as representations while dropping their ad-hoc implementational constraints, allowing us to consider a much larger solution space while still keeping compact recursive numeral representations. Another complementary path would be to provide a mechanistic explanation of how these systems have been shaped over time via general learning mechanisms such as reinforcement learning, similar to carlsson2021learning and silvi2025le, and also via behavioural experiments.